steve is 4 years older than rose. In 5 years from now, the sum of their ages will be equal to twice the sum of their ages from 5 years ago. what is steve's present age
Answer by TimothyLamb(4379)

x = steve's age now
y = rose's age now
---
x = y + 4
(x + 5) + (y + 5) = 2((x - 5) + (y - 5))
x + 5 + y + 5 = 2(x - 5 + y - 5)
x + 5 + y + 5 = 2x - 10 + 2y - 10
x + y + 10 = 2x + 2y - 20
x + y = 30
---
x = y + 4
x + y = 30
---
put the system of linear equations into standard form
---
x - y = 4
x + y = 30
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = steve's age now = 17
y = rose's age now = 13
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php